# Enabling Physical Localization of Uncooperative Cellular Devices

### v2

@@ Abstract
In cellular networks, it can become necessary for authorities to physically locate user devices for tracking criminals or illegal devices. While cellular operators can provide authorities with cell information the device is camping on, fine-grained localization is still required. Therefore, the authorized agents trace the device by monitoring its uplink signals. However, tracking the uplink signal source without its cooperation is challenging even for operators and authorities. Particularly, three challenges remain for fine-grained localization: i) localization works only if devices generate enough uplink traffic reliably over time, ii) the target device might generate its uplink traffic with significantly low power, and iii) cellular repeater may add too much noise to true uplink signals. While these challenges present practical hurdles for localization, they have been overlooked in prior works.
@@
In this work, we investigate the impact of these real-world challenges on cellular localization and propose an Uncooperative Multiangulation Attack (UMA) that addresses these challenges. UMA can 1) force a target device to transmit traffic continuously, 2) boost the target’s signal strength to the maximum, and 3) uniquely distinguish traffic from the target and the repeaters. Notably, the UMA technique works without privilege on cellular operators or user devices, which makes it operate on any LTE network. Our evaluations show that UMA effectively resolves the challenges in real-world environments when devices are not cooperative for localization. Our approach exploits the current cellular design vulnerabilities, which we have responsibly disclosed to GSMA.
1 Introduction
@@
The inherent broadcast nature of wireless signals enables seamless communication without the need for revealing precise physical locations. This very feature, however, poses a significant challenge for authorities with legitimate reasons to pinpoint the exact location of targeted user devices. For example, the perpetrators using SIM boxes engaged in activities like bypass fraud [42, 44] and scams [15, 26] can operate in densely populated urban environments, making it difficult for law enforcement agencies to track and apprehend them. In cases involving criminal activities, it becomes imperative to locate and arrest individuals promptly based on the signals emitted by their cellular devices.
@@
Authorities may seek assistance from mobile operators in the process of cellular device localization, and operators are generally willing to cooperate if the requests are valid and comply with legal procedures [18]. However, mobile operators’ assistance in localization comes with certain limitations. While they can provide authorities with readily-available, high-level user data, such as serving cell ID or triangulation based on signal strength, the level of precision achieved remains relatively coarse-grained. For example, Cell ID (CID) provides only limited positioning of devices with about 1km accuracy [13, 17, 41]. While existing native 3GPP positioning features, such as Observed Time Difference of Arrival (OTDoA) via Positioning Reference Signal (PRS), Uplink Time Difference of Arrival (UTDoA), E-CID (Enhanced CID), have been proposed to improve cellular localization, they are all optional features [3] and thus cannot provide universal fine-grained localization across all mobile operators, radio-access network (RAN) vendors, and User Equipment (UE) manufacturers.
@@
For fine-grained LTE localization localizers require active, real-time monitoring of UE uplink signals, the true signal sources of targeted UEs. This requires additional physical-layer approaches beyond collecting control-plane data in mobile networks; like prior works that follow. For example, Timing Advance (TA) commands used in cellular networks account for the propagation delay of radio signals between the base station and a UE and it can be used to infer sub-cell-level UE locations [48]. Localizers may also determine the distance to a target user by utilizing Time of Arrival (ToA) measurements; e.g., LTrack [31] exploits the ToA of UE uplink signals and the TA commands in the LTE systems. LTEye [32] measures the Angle of Arrival (AoA) of UE uplink signals using a Synthetic Aperture Radar (SAR).
Unfortunately, however, whether these physical-layer cellular localization techniques can provide reliable, fine-grained localization of cellular devices in practice has not been thoroughly tested and demonstrated. Thus, the effectiveness of these techniques under realistic scenarios within a universal cellular infrastructure remains unclear. As a result, this work aims to a) reveal any overlooked challenges of cellular localization in real-world environments, b) address them in a universal setting, and c) demonstrate an end-to-end fine-grained cellular localization.
@@
In this paper, we first investigate and summarize three real-world hurdles in applying known physical-layer localization techniques on UE uplink signals through real-world passive experiments. C1) First, a localizer is able to track a UE only when it generates enough uplink traffic reliably for a certain time period. This means that if a target UE is not actively transmitting signals, the localizer cannot effectively locate its uplink signals. C2) Second, when a target is in close proximity to the cell tower, its uplink signal power may be significantly reduced. This poses a challenge for localizers, as detecting and accurately determining the direction of the target’s uplink signal becomes challenging in urban cellular networks where base stations are densely deployed. C3) Last, cellular repeaters, which are ubiquitous in modern indoor environments, amplify and relay downlink/uplink signals and this introduces severe noise to the localization of true UE uplink signals.
@@
It is noteworthy that cellular operators’ assistance is not sufficient to address the above challenges. Cellular operators have full control over their RAN infrastructure, including base stations (eNBs), repeaters, and UEs; however, their control is limited when it comes to the uplink signals of specific target UEs. RAN vendors offer many management features to mobile operators via proprietary interfaces, allowing them to operate the RAN infrastructure. Nevertheless, these do not extend to making arbitrary changes to the uplink signals of a particular UE. To the best of the authors’ knowledge, no RAN systems exist that allow mobile operators to finely adjust uplink scheduling or power-control, which are necessary to address the challenges of achieving reliable fine-grained cellular localization.
@@
Real-world case: Vishing investigation. As a realistic example, police agents in certain countries trace the location of the cellular devices engaged in vishing fraud called SIM Boxes [22, 23, 44], and subsequently confiscate them [38, 45]. Note that those SIM Boxes are uncooperative devices. They don’t have GPS sensors and only have call-relaying capability. The agents are initially provided with SIM boxes’ coarse-grained location (i.e., cell ID) from mobile operators. They then employ a portable sniffer to track uplink signals emitted by SIM boxes. In this situation, they may encounter the aforementioned challenges, especially in urban cellular networks, and cellular operator’s assistance is insufficient to address them.
@@
To overcome these hurdles, we design an Uncooperative Multiangulation Attack (UMA) and show that overcoming these challenges is indeed feasible with UMA. We show that a localizer with UMA, after obtaining coarse-grained location information of a target UE from the mobile operator, can further obtain UE’s strong uplink signals reliably. We show via end-to-end demonstration in this line of work that real-time manipulation of downlink and uplink signals is feasible and thus physical-layer based uncooperative localization becomes possible with the current, limited operator support. As it requires minimal operator support (e.g., only cell-level information) and works in any LTE networks that follow the mandatory standard features, UMA achieves not only reliable but also universal cellular localization.
@@
(2) We propose a power boosting, designed to resolve C2 and C3. Using power boosting, the localizer, now, impersonates the eNB and instructs the target UE to increase its transmission power to the maximum level. The power boosting directly addresses C2, as it allows the localizer to increase the target’s signal to its maximum level. Moreover, the power boosting enables the localizer to overcome the impact of a repeater. Since the repeater consistently transmits with constant signal strength, executing the power boosting allows the localizer to observe a notable disparity in signal strength between the target UE and the repeater. Hence, this approach effectively resolves challenge C3 as well.
@@
  • Universal. While existing native 3GPP positioning (e.g., OTDOA via PRS, E-CID) depends on optional features in UE implementations [3], UMA operates effectively in virtually any LTE networks that adhere to the core, mandatory standard features.
@@
  • End-to-end. UMA provides publicly available end-to-end demonstration (e.g., from a target’s phone number to its physical location tracking) of a fine-grained cellular localization technique; see our public demonstration video: https://www.youtube.com/watch?v=Y0MJNQCKs7I.
@@
Note that localization techniques are often regarded as privacy attacks as they could violate user privacy depending on the context in which the techniques are used. While we focus on scenarios that assume authorities with legitimate purposes as the main performer of the localization techniques, we acknowledge that unprivileged third parties may use the UMA technique for unlawful purposes. To simplify writing and our arguments, in this paper, we decide to refer the UMA technique as attacks.
@@
Responsible disclosure. We have reported our findings to GSMA. They have acknowledged our contributions with CVD–2023-70 and CVD–2023-77.
2 Preliminaries
In this section, we provide the technical backgrounds of LTE necessary to understand uncooperative localization.
@@ 2 Preliminaries
Identifiers in LTE. When a UE communicates with LTE networks, several identifiers are involved, which are assigned by different network entities. The identifiers can be classified as permanent and temporary. Representatively, the International Mobile Subscriber Identity (IMSI) is a permanent identifier that remains unchanged once assigned to the user’s User Identity Module (USIM).
@@
During radio communication, two temporary identifiers – Globally Unique Temporary Identifier (GUTI) and Radio Network Temporary Identifier (RNTI) – are mainly used to avoid privacy issues [20, 29] caused by the permanent identifiers. GUTI is assigned by core networks when a UE accesses LTE cellular networks and a connection is established. RNTI is assigned by an eNB for physical layer communication in cellular radio networks when the UE requests a connection to radio access networks [4].
RNTI changes periodically whenever a radio connection is re-established between a UE and an eNB. For example, when there is no data transmission between the UE and the eNB for a specific period, the eNB removes the connection for network resource and UE’s battery management. Then, when the UE needs data transmission, a new RNTI is allocated to the UE.
Figure 1: Physical layer channel structure.
Downlink Control Information (DCI). As shown in Figure 1, the eNB and UE transmit their traffic through downlink and uplink channels in the physical layer, respectively. In the physical layer channel, the subframe is a fundamental resource unit for radio communication.
@@
In order to decode downlink data and enable the transmission of uplink data, Downlink Control Information (DCI) plays a crucial role, especially in notifying scheduling information and transmission power control for the UEs. Each UE utilizes the RNTI to decode the DCI message and identify the allocated resources. Specifically, a certain format of DCI, DCI format 0 (DCI 0), is used for controlling uplink transmission power to ensure optimal signal quality and identifying allocated resources (i.e., uplink grant) on Physical Uplink Shared Channel (PUSCH). Note that DCI 0 is broadcast without encryption, it is possible to obtain a UE’s uplink scheduling information by using the UE’s RNTI. The process of DCI decoding and resource allocation is illustrated in Figure 1.
3 Attack Model
3.1 Motivation
@@
Beyond-cell-level, fine-grained localization of cellular devices has been a highly desired feature for various applications, such as emergency response, criminal investigation, and even unlawful stalking or surveillance. Localizers begin by obtaining the coarse-grained location of a target device (e.g., directly obtaining its serving cell ID from the mobile operator [18] or inferring the serving cell via paging side channel [46]), and then attempt to further narrow down.
@@
However, existing studies show that fine-grained cellular localization is hard even for authorized agents with assistance from mobile operators. A survey from Electronic Frontier Foundation in 2021 [21] shows that mobile signal tracking via triangulation by operators offer only about 1 km accuracy. It also discusses other techniques, such as cell site simulators, which are used by law enforcement agencies; however, they allow only coarse-grained device presence detection, and works only when the target device supports 2G, a soon-to-be-obsolete technology in the market. Peral-Rosado et al[17] also have surveyed localization technologies that can be used by mobile operators. It discusses a few technologies that have potential to provide beyond-cell-level localization (e.g., 3GPP’s E-CID, OTDOA, and UTDOA) but they are only optional features [3] and thus cannot provide universal fine-grained localization across all mobile operators, RAN vendors, and UE manufacturers. As an aside, GPS information can be a useful means for authorities to trace the location of the target UE. However, authorities cannot access GPS location without the permission of mobile users when they are using iPhone [36] and devices employed by criminals may lack support for GPS features.
Figure 2: Idealized view vs. real-world view.
@@
Idealized view on cellular localization. Understanding the limited cellular localization capability in modern cellular networks, there have been several academic studies that attempt to achieve fine-grained localization of cellular devices outside the official localization features provided by mobile operators [31, 32]. These studies, however, have rather idealized views on the cellular localization problem, thus overlooking several practical challenges that a localizer may suffer when attempting physical localization in real-world environments.  Figure 2 illustrates this idealized view on cellular localization used in previous studies. In this view, a target device is always transmitting uplink signals, which are always strong enough to be detected by the localizer. Also, strictly only one uplink signal is transmitted by the target device at a time. Thus, the localizer can conveniently conduct physical-layer localization techniques (e.g., AoA, ToA, TA) to locate the target’s uplink signal source, after they obtain the target device’s temporary network identifier (e.g., RNTI) and detect its uplink signal.
@@
While this simplifed view is a good starting point to understand the cellular localization problem, in a more realistic, network-wide view on cellular localization, the localizer immediately faces several challenges because the real RAN environment is much more complex. First, targeted UEs may not be transmitting uplink signals at the time the localizer is attempting to trace them. This not only makes localization hard but also makes it difficult to identify the target UE’s temporary network identifier. Second, target UE’s uplink signal power varies significantly because it is always adjusted depending on its channel condition. Last, the localizer sees not only the target UE’s uplink signal but also many replicas of the same signal amplified and relayed by repeaters that are widely deployed to enhance cellular channel quality.
3.2 Threat Model

### v3

@@ Abstract.
In cellular networks, authorities may need to physically locate user devices to track criminals or illegal equipment. This process involves authorized agents tracing devices by monitoring uplink signals with cellular operator assistance. However, tracking uncooperative uplink signal sources remains challenging, even for operators and authorities. Three key challenges persist for fine-grained localization: i) devices must generate sufficient, consistent uplink traffic over time, ii) target devices may transmit uplink signals at very low power, and iii) signals from cellular repeaters may hinder localization of the target device. While these challenges pose significant practical obstacles to localization, they have been largely overlooked in existing research.
@@
This work examines the impact of these real-world challenges on cellular localization and introduces the Uncooperative Multiangulation Attack (UMA) to address them. UMA can 1) force a target device to transmit traffic continuously, 2) boost the target’s signal strength to maximum levels, and 3) uniquely differentiate between signals from the target and repeaters. Importantly, UMA operates without requiring privileged access to cellular operators or user devices, making it applicable to any LTE network. Our evaluations demonstrate that UMA effectively overcomes practical challenges in physical localization when devices are uncooperative.
Cellular localization; LTE; Law enforcement
journalyear: 2024copyright: rightsretainedconference: The 30th Annual International Conference on Mobile Computing and Networking; November 18–22, 2024; Washington D.C., DC, USAbooktitle: The 30th Annual International Conference on Mobile Computing and Networking (ACM MobiCom ’24), November 18–22, 2024, Washington D.C., DC, USAdoi: 10.1145/3636534.3690709isbn: 979-8-4007-0489-5/24/11ccs: Networks Mobile and wireless security
1. Introduction
@@
The broadcast nature of wireless signals allows seamless communication without revealing precise locations. However, this feature challenges authorities who need to pinpoint specific user devices for legitimate reasons. For example, the perpetrators engaged in criminal activities (Reaves et al., 2015; Sahin et al., 2017; Corporation, 2021; Houweling, 2021) can operate their cellular devices (e.g., SIM boxes) in densely populated urban environments, making it difficult for law enforcement agencies to track and apprehend them.
@@
Authorities may seek assistance from mobile network operators (MNOs) in the process of cellular device localization, and operators are generally willing to cooperate if the requests are valid and comply with legal procedures (Downing, 2016). However, MNOs’ assistance in localization comes with certain limitations. While they can provide authorities with readily-available, high-level user data, such as serving cell ID or triangulation based on signal strength, the level of precision achieved remains relatively coarse-grained. For example, Cell ID (CID) provides only limited positioning of devices with about 1km accuracy (Cherian and Rudrapatna, 2013; del Peral-Rosado et al., 2017; Karri Ranta-Aho, 2010). While existing native 3GPP positioning features, such as Observed Time Difference of Arrival (OTDoA), Uplink Time Difference of Arrival (UTDoA), Enhanced CID (E-CID), have been proposed to improve cellular localization, they are all optional features (3GPP. TS 36.305, v17.0.0, 2022) and thus cannot provide universal fine-grained localization across all MNOs, radio-access network (RAN) vendors, and User Equipment (UE) manufacturers.
@@
Fine-grained LTE localization requires active, real-time monitoring of uplink signals from targeted UEs. This necessitates advanced physical-layer approaches beyond conventional control-plane data collection in mobile networks. For instance, Timing Advance (TA) commands, which compensate for signal propagation delays between base stations and UEs, can be used to infer sub-cell-level UE locations (Tan et al., 2021). Localizers may also employ Time of Arrival (ToA) measurements to determine a target user’s distance. LTrack (Kotuliak et al., 2022) leverages both ToA of UE uplink signals and TA commands, and LTEye (Kumar et al., 2014) measures the Angle of Arrival (AoA) of UE uplink signals using Synthetic Aperture Radar (SAR).
While existing physical-layer approaches show promise, their effectiveness for reliable fine-grained localization of uncooperative devices – those that are neither directly controlled by the localizer nor report their own location – in real-world scenarios remains unproven. The gap between their performance in controlled environments and our goal of localizing uncooperative devices in realistic scenarios poses significant challenges. This work aims to a) identify real-world obstacles overlooked in cellular localization, b) develop universally applicable solutions, and c) demonstrate a comprehensive process for fine-grained localization of uncooperative cellular devices.
@@
In this paper, we conduct real-world experiments to investigate challenges in applying existing physical localization techniques to uplink signals of uncooperative devices. Our experiments reveal three key issues: C1) Insufficient uplink traffic: Localization becomes impossible if a target UE is not actively transmitting uplink traffic during tracking. C2) Low signal power: When a target is close to a cell tower, its uplink signal power may be significantly reduced, hindering signal detection and location determination, especially in urban areas with densely deployed base stations. C3) Interference from repeaters: Cellular repeaters, ubiquitous in modern indoor environments, amplify and relay cellular signals, introducing severe noise in determining a target’s location.
@@
It’s important to note that cellular operators’ assistance alone is insufficient to address these challenges. While operators have full control over their RAN infrastructure, including base stations (eNBs) and repeaters, and RAN vendors offer various management features via proprietary interfaces, these capabilities do not extend to making arbitrary adjustments to specific UEs’ uplink signals. To the best of our knowledge, no existing RAN systems allow operators to finely tune uplink scheduling or power control of a particular UE, which are necessary for achieving reliable, fine-grained cellular localization.
@@
Real-world case: Vishing investigation. As a realistic example, police agents in East Asia trace the location of the cellular devices engaged in vishing fraud called SIM boxes (GSMA, 2020, 2024; Sahin et al., 2017), and subsequently confiscate them (Chae, 2022). Note that those SIM Boxes are uncooperative devices. They don’t have GPS sensors and only have call-relaying capability. The agents are initially provided with SIM boxes’ coarse-grained location (i.e., cell ID) from MNOs. They then employ a portable sniffer to track uplink signals emitted by SIM boxes. In this situation, they may encounter the aforementioned challenges, especially in urban cellular networks, and cellular operator’s assistance is insufficient to address them.
@@
To overcome these hurdles, we design an Uncooperative Multiangulation Attack (UMA) and show that overcoming these challenges is indeed feasible with UMA. We reveals that a localizer with UMA, after obtaining coarse-grained location information of a target UE from the MNO, can further obtain UE’s strong uplink signals reliably. We demonstrate end-to-end that real-time manipulation of downlink and uplink signals is feasible, enabling physical localization of uncooperative devices with current, limited operator support. As UMA functions in any LTE network adhering to mandatory standard features, it achieves not only reliable but also universal cellular localization.
@@
(2) We propose a power boosting, designed to resolve C2 and C3. Using power boosting, the localizer, now, impersonates the eNB and instructs the target UE to increase its transmission power to the maximum level. The power boosting directly addresses C2, as it allows the localizer to increase the target’s signal to its maximum level. Moreover, the power boosting enables the localizer to overcome the impact of a repeater. Since the repeater consistently transmits with constant signal strength, executing the power boosting allows the localizer to observe a notable disparity in signal strength between the target UE and the repeater.
@@
  • Universal. While existing native 3GPP positioning depends on optional features in eNB/UE implementations (3GPP. TS 36.305, v17.0.0, 2022), UMA operates effectively in virtually any LTE networks that adhere to the core, mandatory standard features.
@@
  • End-to-end. UMA provides publicly available end-to-end demonstration (i.e., from a target’s phone number to its physical location) of fine-grained cellular localization; see our public website for its demonstration videos (UMA, 2024).
@@
Note that localization techniques are often regarded as privacy attacks as they could violate user privacy depending on the context in which the techniques are used. While we focus on scenarios that assume authorities with legitimate purposes as main performers of localization techniques, we acknowledge that unprivileged third parties may use UMA for unlawful purposes. To simplify writing and our arguments, in this paper, we refer to the UMA technique as attacks.
@@
Responsible disclosure. We have reported our findings to GSMA. They have acknowledged our contributions with CVD-2023-0070 and CVD-2023-0077.
2. Preliminaries
@@
Identifiers in LTE. When a UE communicates with LTE networks, several identifiers are involved, which are assigned by different network entities. The identifiers can be classified as permanent and temporary. Representatively, the International Mobile Subscriber Identity (IMSI) is a permanent identifier that remains unchanged once assigned to the user’s Universal Subscriber Identity Module (USIM).
@@
During radio communication, two temporary identifiers – Temporary Mobile Subscriber Identity (TMSI) and Radio Network Temporary Identifier (RNTI) – are mainly used to avoid privacy issues (Cheshire, 2015; Goodin, 2013) caused by the permanent identifiers. TMSI is assigned by core networks when a UE accesses LTE cellular networks and a connection is established. RNTI is assigned by an eNB for physical layer communication in cellular radio networks when the UE requests a connection to radio access networks (3GPP. TS 36.321, v16.6.0, 2021). RNTI changes occur when a UE – eNB radio connection is re-established. If no data transmission occurs for a specified period, the eNB terminates the connection. A new RNTI is then allocated to the UE for its next data transmission.
@@
Downlink Control Information (DCI). The eNB and UE transmit their traffic through downlink and uplink channels in the physical layer, respectively. In the physical layer channel, the subframe is a fundamental resource unit for radio communication. In order to decode downlink data and enable the transmission of uplink data, DCI plays a crucial role, especially in notifying scheduling information and transmission power control for the UEs. Each UE utilizes the RNTI to decode the DCI message and identify the allocated resources. Specifically, a certain format of DCI, DCI format 0 (DCI 0), is used for controlling uplink transmission power to ensure optimal signal quality and identifying allocated resources (i.e., uplink grant) on Physical Uplink Shared Channel (PUSCH). Note that DCI 0 is broadcast without encryption, it is possible for a third party to obtain a specific UE’s uplink scheduling information by using the UE’s RNTI.
3. Attack Model
3.1. Motivation
@@
Fine-grained localization of cellular devices beyond cell-level precision is highly desired for various applications, including emergency response, criminal investigation, and potentially unlawful surveillance. Localizers typically start with a device’s coarse-grained location before attempting to narrow it down further.
@@
However, studies indicate that even authorized agents with mobile operator assistance struggle to achieve fine-grained cellular localization. While authorities can obtain approximate location from operators, this information alone is insufficient for localizing cellular devices “down to the front door,” especially when dealing with uncooperative devices. A survey from Electronic Frontier Foundation (Foundation), 2021) shows that mobile signal tracking via triangulation by operators offers only about 1 km accuracy. It also discusses other techniques, such as cell site simulators, which law enforcement agencies use; however, they allow only coarse-grained device presence detection, and work only when the target device supports 2G, a soon-to-be-obsolete technology in the market. Peral-Rosado et al(del Peral-Rosado et al., 2017) also have surveyed localization technologies that can be used by mobile operators. It discusses a few technologies that have potential to provide beyond-cell-level localization (e.g., E-CID, OTDoA, and UTDoA) but they are optional features (3GPP. TS 36.305, v17.0.0, 2022) and thus cannot provide universal fine-grained localization across all MNOs, RAN vendors, and UE manufacturers. As an aside, GPS information can be a useful means to trace the target UE. However, authorities cannot access GPS location without the permission of mobile users when they are using iPhone (McConnell, 2017) and devices employed by criminals may lack support for GPS features.
Figure 1. Idealized view vs. real-world view.
@@
Idealized view on cellular localization. There have been several academic studies that attempt to achieve fine-grained localization of cellular devices outside the official localization features provided by mobile operators (Kotuliak et al., 2022; Kumar et al., 2014). These studies, however, have rather idealized views on the cellular localization problem, thus overlooking several practical challenges that a localizer may suffer when attempting physical localization in real-world environments.  Figure 1 illustrates this idealized view on cellular localization used in previous studies. In this view, a target device is always transmitting uplink signals, which are always strong enough to be detected by the localizer. Also, strictly only one uplink signal is transmitted by the target device at a time. Thus, the localizer can conveniently conduct physical-layer localization techniques (e.g., AoA, ToA) to locate the target’s uplink signal source, after they obtain the target device’s temporary network identifier (e.g., RNTI) and detect its uplink signal.
@@
While this simplified view is a good starting point to understand the cellular localization problem, in a more realistic, network-wide view on cellular localization, the localizer immediately faces several challenges because the real RAN environment is much more complex. First, targeted UEs may not be transmitting uplink signals at the time the localizer is attempting to trace them. This not only makes localization hard but also makes it difficult to identify the target UE’s temporary network identifier. Second, target UE’s uplink signal power varies significantly because it is always adjusted depending on its channel condition, which may prevent the localizer from detecting low power signals. Last, the localizer sees not only the target UE’s uplink signal but also many replicas of the same signal amplified and relayed by repeaters that are widely deployed to enhance cellular channel quality.
3.2. Threat Model